# A Survey on Data Selection for Language Models

### v2

@@ 3.4 Domain-Specific Selection
While some language models are pretrained for general domain use, they can also be trained with a specific domain in mind, where selection can assist in finding data similar to the desired domain’s distribution (e.g. medicine or law). Domain-specific filtering methods mostly assume access to some in-domain data and additional data from auxiliary sources. The goal of filtering is then to find the auxiliary data which most resembles the distribution of in-domain data.
@@
Engstrom et al. (2024) propose a model-dependent alternative to Moore-Lewis-based methods, which only depend on the data. They propose to use datamodels, a utility function that can approximately map a subset of training data to model performance after training on that subset. Generally, estimating this requires training some models with different subsets of data. At very high filtering ratios (less than 25% of original dataset size), they demonstrate improved performance over DSIR (Xie et al., 2023b), but require significantly more compute.
In general, methods derived from Moore-Lewis selection have been the gold-standard, but it may be possible to adopt other methods of data attribution and valuation, similar to datamodels, for the purpose of selecting in-domain data. However, these methods typically have higher computational requirements, a challenge that needs to be addressed. Perplexity-based methods, similar to perplexity-based quality filtering methods (§3.3), which only require an in-domain model are also a plausible research direction.
@@
Prior methods of domain-specific selection utilize just a model of the in-domain data, but these methods can skew the resultant distribution towards the modes of the existing in-domain data (Sethy et al., 2006). For example, it is possible that the in-domain data contains some very common phrases (e.g. “okay”) which are not necessarily representative of the domain. To avoid this issue, it is possible to perform deduplication on the in-domain data. However, a delicate balance must be reached, as deduplication can skew the in-domain data away from valid modes. Domain-specific selection methods that utilize a general-domain model in addition to the in-domain model (e.g. Moore-Lewis selection) handle this issue quite well and thus it is highly recommended to use such methods.
Domain-specific selection aims to find data most similar to a specific domain (e.g. medicine or law). Moore-Lewis style methods (requiring one in-domain model and one general purpose model) have stood the test of time and are still the most popular methods for domain-specific filtering. [Uncaptioned image] Datamodels and other methods based on data attribution and valuation can be explored as competitive methods, but generally have higher computational requirements. [Uncaptioned image] Perplexity-based methods may be another direction to explore. They require only an in-domain model, which could improve efficiency. It’s possible that they would not perform as well as Moore-Lewis based methods, but because they use less compute they could be used in multiple rounds of selection, as in bootstrapping. Domain-Specific Selection Summary & Ideas
@@ 3.5 Data Deduplication
Datasets originating from internet dumps often abound with duplicate and near duplicate documents (Elazar et al., 2023; Magnusson et al., 2023). A dataset with duplicated data, or near duplicates, increases the distribution density around those areas. In cases where these high-density points are of high value, keeping near duplicates can be beneficial. However, for pretraining where the exact evaluation distribution is unknown, it is generally preferred to remove duplicates so that the training distribution provides greater coverage with less redundancy. Carlini et al. (2023) demonstrate a direct relation between the number of times a data point is duplicated and the degree to which a model will memorize the data point, a relation which increases with model scale. Additionally, filtering out duplicates can reduce the training time of machine learning models (Sorscher et al., 2022) and has been shown to sometimes improve accuracy on downstream tasks (Lee et al., 2022a; Tirumala et al., 2024).
@@
Some exact-matching utility functions utilize hash functions. These hash-based methods are guaranteed to find all exact matches but, due to the possibility of collisions, may also find false positives (accidentally removing non-matching documents), although Elazar et al. (2023) demonstrate that a simple hash function on large scale datasets can experience no collisions. CCNet (Wenzek et al., 2020) performs deduplication by first normalizing all text, then computing a hash code for each document, and finally comparing the first 64-bits of SHA-1 digits of the hashed documents. When developing OSCAROrtiz Suárez et al. (2019) use a simple non-collision resistant hashing algorithm to determine whether whole documents are exact matches, and. Another common method for deduplication is the Bloom filter which utilizes a space-efficient bit array to compare hashed documents (Bloom, 1970; Soldaini et al., 2024).
@@
Similarly to exact matching methods, hash-based methods are very commonly used for approximate matching methods. For example, MinHash (Broder, 1997) is an approximate hashing algorithm that excels at finding templated documents (e.g. licenses with only specific entities differing or placeholder SEO text repeated across websites). Modern deduplication methods use other variants of hashing such as MinHashLSH (Brown et al., 2020) and SimHash (Charikar, 2002) that utilize locality sensitive hashing. [^11] See the blogpost by Mou (2023) for implementation details on MinHashLSH The general method behind hashing methods for deduplication is to first group small segments of each documents (often called shingling), then to encode each segment using a large number of hashing functions into features. Next, aggregate the features (e.g. MinHashLSH keeps only the smallest features). Finally, documents can be compared with some method of similarity computation between features (e.g. MinHashLSH uses a Jaccard similarity) and removing documents below some threshold. When creating MassiveTextRae et al. (2021) use MinHash with 13-grams and 450 hash functions and set the jaccard similarity threshold at 0.8, randomly removing either of the documents that are determined to be duplicates. In contrast, Gao et al. (2020) use only 10 hash functions, but Rae et al. (2021) find that the more aggressive deduplication leads to better model performance.
@@
One concern with hashing-based method for approximate deduplication is that because hashing functions do not consider the order of each individual shingle, they act similarly to bag-of-word models where long documents are more likely to have similar representations. Laurençon et al. (2022) found that this led to long documents having a higher percentage of false positive as duplicates. Thus, they did not remove documents in a cluster of near-duplicates if a document is longer than 6000 characters. Instead, they use substring deduplication based on the suffix array (as in Lee et al. (2022a)) for documents with more than 6000 characters.
Methods can search for duplicated documents, or just chunks of text which are duplicated within and across documents. Once duplicated chunks have been found, the question arises: what to do with these chunks? In the ExactSubstr method, Lee et al. (2022a) proposed to remove only the exact substring that is duplicated and Penedo et al. (2023) also elect to remove only the duplicated span and keep the rest of the document. This can lead to incoherent documents, but Penedo et al. (2023) empirically find it to work better than removing the entire document. Another option is to loss-mask the duplicated tokens (prevent the loss from being computed on duplicated text), but this adds code complexity as the start and end of each duplicated span will need to be tracked and considered when training the model.
Deduplication is a very important step and is often used multiple times within the data selection pipeline. First with a simple URL-based filter, then with more complicated hashing- and model-based methods. Deduplication generally relies one of four methods: URLs, hashing, string metrics, or model representations. Bloom filters are often used as state-of-the-art exact deduplication (even though they can remove non-duplicated data) as they are very space efficient. For a more thorough deduplication, approximate matching methods can be used, where MinHashLSH is currently the most commonly used method. [Uncaptioned image] In settings where data efficiency is the primary goal, model-based deduplication has demonstrated good results. However, these methods are relatively new and need further study to demonstrate their benefits in other cases (e.g. when performance is the primary goal). Data Deduplication Summary & Ideas
@@ 3.6 Filtering Toxic and Explicit Content
Practitioners often aim to curate a dataset that is “unbiased and fair”. As such, the aim is for models trained on such data to follow similarly unbiased behavior. Certain popular datasets have also been found to have illegal or extremely inappropriate content, which motivates strict data filters (David, 2023). Towards this purpose, most datasets employ a process to filter out documents that do not conform with such behaviors. Bias and fairness are one of the least well-defined filtering requirements and may be in conflict with different ideologies (Gururangan et al., 2022) and implications (Longpre et al., 2023c). A crucial aspect of this topic is the lack of transparency (Bommasani et al., 2023), requirements, and the effects of such bespoke bias filtering from pretraining data on downstream behavior. Many works solely provide minimal information about the heuristics used for filtering out “biased” data without reporting any ablation on the impact on downstream tasks. Longpre et al. (2023c) being one of the only papers that controls and ablates for some of these filters to comprehensively study their effects. They find that using toxicity filters (specifically, Jigsaw’s Perspective API [^12] https://www.perspectiveapi.com/) leads to models that produce less toxic texts, however, the model’s ability to detect toxic content also decreases. As such, depending on the end-application one must be careful when using such filters.
In what follows we summarize the filters related to bias and fairness used in different papers based on the filter types. As mentioned above, such filters are often used due to some heuristics and the effect of such filtering is not well studied.
@@
Much of the existing NSFW content can be filtered solely based on the source URL, a typical metadata stored alongside the text of datasets. Datasets such as RefinedWeb (Penedo et al., 2023) filter out documents coming from URLs that match words from a curated list, as well as from blacklisted domains.
@@
A less common method is to score documents using a model trained only on the undesirable content. Jansen et al. (2022) train a model directly on adult and harmful content, and then evaluate the model on documents, using the perplexity as a utility function where documents with a high perplexity are assumed to not contain any adult or harmful content. Of course, there are limitations to this method. It is important to train the harmful content model on a wide variety of harmful content, otherwise this method may struggle to detect new types of harmful content.
@@
Finally, Personally Identifiable Information (PII) is another source of concern, which may expose the personal information of individuals. Frequent targeted PIIs are phone numbers, email addresses, IP addresses, and secret keys. A common approach for handling PII is to detect them using regex (Subramani et al., 2023; Elazar et al., 2023; Soldaini et al., 2024; Groeneveld et al., 2024) or classifiers (Allal et al., 2023). Once PII terms are detected, a common practice is to obfuscate them with special tokens.
There is a wide variety of toxic and explicit content that are generally undesirable to have in training or evaluation data. Common methods for filtering out toxic and explicit content include blacklisting certain web domains, blacklisting certain words, using a classifier to detect harmful content, using perplexity to detect harmful content, and removing PII with regular expressions or classifiers. [Uncaptioned image] There have been limited studies on the effects of these filters, warranting further research on the impacts of such filtering to understand how they affect the distribution of data and the downstream models behaviors. [Uncaptioned image] A useful direction for research is how to mitigate any negative impacts of this filtering. Or conversely, how can toxic and explicit behaviors be reduced at other stages of model development (e.g. alignment), or possibly even at evaluation time through decoding methods. Toxic and Explicit Content Filtering Summary & Ideas
@@ 3.7 Specialized Selection for Multilingual Models
Many selection methods, such as deduplication, can be reused across languages (Laurençon et al., 2022). However, it is critical to adapt the hyperparameters for the respective language and often manually tune them (Laurençon et al., 2022; Scao, 2023; Scao et al., 2022; Workshop et al., 2022). For example, while heuristics based on technical characters ([0-9{}+/()>]) are still valid, for a length-based filter (discussed in §3.2) one needs to consider that languages like Chinese have much higher information density per character leading to the same content requiring fewer characters than English. Thus, in a filtering pipeline Chinese would require a smaller cut-off value for minimum text length than English (Scao, 2023). Additionally, when using language identification models (e.g. fastText) to detect the language of a document, it is important that the threshold for removal is determined by native speakers, and specific to each language (Laurençon et al., 2022). Not only is the language ID used for determining the language of a document, but it can also be used at the sentence level, where Kudugunta et al. (2023) consider removing sentences if the sentence level language ID and document level ID do not match.
@@
Languages with non-Latin scripts (e.g. Mandarin, Arabic, and Hindi) commonly suffer from incorrect encoding which may require additional filtering or noise correction (Kudugunta et al., 2023). Further, some scripts such as Dongba symbols are not yet supported by Unicode, making the selection of such data especially difficult. Using script-based filters can also serve as language identification for languages that have a one-to-one mapping with a script, such as Japanese Hiragana (§3.1).
@@
Many of the selection methods from other sections in this work can simply be reused when filtering for multilingual data (e.g. deduplication, heuristics on technical characters, data mixing), but some require special parameters for each language (e.g. length-based heuristics). [Uncaptioned image] Particularly for low-resource languages, native speakers must be incorporated into the filtering process to ensure that the data maintains an acceptable quality. Aya (Singh et al., 2024) is a great example of how this type of inclusion and collaboration can look. Specialized Selection for Multilingual Models Summary & Ideas
@@ 3.8 Data Mixing
Large pretraining corpora are often comprised of a mixture of data from $k$ domains, where domains are typically defined by provenance. The weighting across domains are specified by the domain weights $\alpha\in\Delta^{k}$, which comprise a $k$-dimensional probability vector. For example, the Pile dataset (Gao et al., 2020) is composed of 24% web data, 9% Wikipedia, 4% GitHub, etc. [^14] The mixture weights depend on the tokenizer. The choice of domain weights can result in significant differences in downstream accuracy (Xie et al., 2023a; Albalak et al., 2023a; Xia et al., 2024a; Fan et al., 2023). Given a fixed training budget, data mixing methods optimize the domain weights to improve training efficiency and model performance.
@@
One challenge that has not been addressed in any of these works is how their weights transfer to new settings. For example, Albalak et al. (2023a) find that the weights determined by Xie et al. (2023a) do not transfer well when using a different tokenizer. It is unclear how to transfer the weights learned under one tokenizer to a new tokenizer without learning the weights from scratch. Furthermore, none of the principled approaches discuss the effects of extrapolating the calculated weights to training on larger quantities of tokens. To exemplify the problem, consider that the weights are calculated by training a model on 50B tokens, and the weights are then used to train a model on 1T tokens. Any domains that saw 1 epoch at 50B tokens, will now see 20 epochs at 1T, leading to possible overfitting on smaller domains.
@@
When a pretraining corpus is composed of data from multiple domains, data mixing is an important step of data selection that assigns weights to each domain, determining the final composition of the dataset. Weights can be determined manually, by heuristics, empirically determined, or from principled approaches. Each of these methods have been proven under specific evaluations, but none has proven as a general-purpose best method. [Uncaptioned image] Data mixing is challenging, as there is a natural trade-off when upweighting one domain, the remaining domains must be downweighted. Research on overcoming this trade-off may be difficult, but would be very impactful. One possible direction are methods that consider the interactions (similarities) between domains, which could identify domains that can be downweighted without reducing performance, and those specific domains whose weights cannot be reduced without detriment to performance. [Uncaptioned image] Each of the approaches discussed has different benefits, and combining approaches (e.g. 1/3 heuristic, 1/3 empirically-drive, 1/3 principled) could lead to better general purpose use, but how to best combine them leads to additional complexity. [Uncaptioned image] Although principled approaches are prominent and have lead to state-of-the-art mixing, the weights they determine can also be brittle when used under different settings (e.g. tokenizer, longer training), which are issues that need to be addressed and explored in future research. Data Mixing Summary & Ideas

### v3

@@ 3.4 Domain-Specific Selection
While some language models are pretrained for general domain use, they can also be trained with a specific domain in mind, where selection can assist in finding data similar to the desired domain’s distribution (e.g., medicine or law). Domain-specific filtering methods mostly assume access to some in-domain data and additional data from auxiliary sources. The goal of filtering is then to find the auxiliary data which most resembles the distribution of in-domain data.
@@
Engstrom et al. (2024) propose a model-dependent alternative to Moore-Lewis-based methods, which only depends on the data. They propose to use datamodels (Ilyas et al., 2022), a utility function that can approximately map a subset of training data to model performance after training on that subset. Generally, producing a datamodel requires training a series of models with different subsets of data. At very high filtering ratios (less than 25% of original dataset size), they demonstrate improved performance over DSIR (Xie et al., 2023b), but require significantly more compute.
In general, methods derived from Moore-Lewis selection have been the gold-standard, but it may be possible to adopt other methods of data attribution and valuation, similar to datamodels, for the purpose of selecting in-domain data. However, these methods typically have higher computational requirements, a challenge that needs to be addressed. Perplexity-based methods (similar to perplexity-based quality filtering methods (§3.3)) that only require an in-domain model are also a plausible research direction.
@@
Prior methods of domain-specific selection utilize just a model of the in-domain data, but these methods can skew the resultant distribution towards the modes of the existing in-domain data (Sethy et al., 2006). For example, it is possible that the in-domain data contains some very common phrases (e.g., “okay”) which are not necessarily representative of the domain. To avoid this issue, it is possible to perform deduplication on the in-domain data. However, a delicate balance must be reached, as deduplication can skew the in-domain data away from valid modes. Domain-specific selection methods that utilize a general-domain model in addition to the in-domain model (e.g., Moore-Lewis selection) handle this issue quite well and thus it is highly recommended to use such methods.
Domain-specific selection aims to find data most similar to a specific domain (e.g., medicine or law). Moore-Lewis style methods (requiring one in-domain model and one general purpose model) have stood the test of time and are still the most popular methods for domain-specific filtering. [Uncaptioned image] Datamodels and other methods based on data attribution and valuation can be explored as competitive methods, but generally have higher computational requirements. [Uncaptioned image] Perplexity-based methods may be another direction to explore. They require only an in-domain model, which could improve efficiency. It’s possible that they would not perform as well as Moore-Lewis based methods, but because they use less compute they could be used in multiple rounds of selection, as in bootstrapping. Domain-Specific Selection Summary & Ideas
@@ 3.5 Data Deduplication
Datasets originating from internet dumps often abound with duplicate and near duplicate documents (Elazar et al., 2024; Magnusson et al., 2023). A dataset with duplicated data, or near duplicates, increases the distribution density around those areas. In cases where these high-density points are of high value, keeping near duplicates can be beneficial. However, for pretraining where the exact evaluation distribution is unknown, it is generally preferred to remove duplicates so that the training distribution provides greater coverage with less redundancy. Carlini et al. (2023) and Kandpal et al. (2022) demonstrate a direct relationship between the number of times a data point is duplicated and the degree to which a model will memorize the data point, a relation which increases with model scale. Additionally, filtering out duplicates can reduce the training time of machine learning models (Sorscher et al., 2022) and has been shown to sometimes improve accuracy on downstream tasks (Lee et al., 2022a; Tirumala et al., 2024).
@@
Some exact-matching utility functions utilize hash functions. These hash-based methods are guaranteed to find all exact matches but, due to the possibility of collisions, may also find false positives (accidentally removing non-matching documents), although Elazar et al. (2024) demonstrate that a simple hash function on large scale datasets can experience no collisions. CCNet (Wenzek et al., 2020) performs deduplication by first normalizing all text, then computing a hash code for each document, and finally comparing the first 64-bits of SHA-1 digits of the hashed documents. When developing OSCAROrtiz Suárez et al. (2019) use a simple non-collision resistant hashing algorithm to determine whether whole documents are exact matches. Another common method for deduplication is the Bloom filter which utilizes a space-efficient bit array to compare hashed documents (Bloom, 1970; Soldaini et al., 2024).
@@
Similarly to exact matching methods, hash-based methods are very commonly used for approximate matching methods. For example, MinHash (Broder, 1997) is an approximate hashing algorithm that excels at finding templated documents (e.g., licenses with only specific entities differing or placeholder SEO text repeated across websites). Modern deduplication methods use other variants of hashing such as MinHashLSH (Brown et al., 2020a) and SimHash (Charikar, 2002) that utilize locality sensitive hashing. [^12] See the blogpost by Mou (2023) for implementation details on MinHashLSH. The general method behind hashing methods for deduplication is to first group small segments of each documents (often called shingling), then to encode each segment using a large number of hashing functions into features. Next, aggregate the features (e.g., MinHashLSH keeps only the smallest features). Finally, documents can be compared with some method of similarity computation between features (e.g., MinHashLSH uses a Jaccard similarity) and removing documents below some threshold. When creating MassiveTextRae et al. (2021) use MinHash with 13-grams and 450 hash functions and set the jaccard similarity threshold at 0.8, randomly removing either of the documents that are determined to be duplicates. In contrast, Gao et al. (2020) use only 10 hash functions, but Rae et al. (2021) find that the more aggressive deduplication leads to better model performance.
@@
One concern with hashing-based method for approximate deduplication is that because hashing functions do not consider the order of each individual shingle (the chunks of text used to create the hashed representation), they act similarly to bag-of-word models where long documents are more likely to have similar representations. Laurençon et al. (2022) found that this led to long documents having a higher percentage of false positive as duplicates. Thus, they did not remove documents in a cluster of near-duplicates if a document is longer than 6000 characters. Instead, they use substring deduplication based on the suffix array (as in Lee et al. (2022a)) for documents with more than 6000 characters.
Methods can search for duplicated documents, or just chunks of text which are duplicated within and across documents. Once duplicated chunks have been found, the question arises: what to do with these chunks? The ExactSubstr method Lee et al. (2022a) removes only the exact substring that is duplicated and Penedo et al. (2023) also elect to remove only the duplicated span and keep the rest of the document. This can lead to incoherent documents, but Penedo et al. (2023) empirically find it to work better than removing the entire document. Another option is to loss-mask the duplicated tokens (prevent the loss from being computed on duplicated text), but this adds code complexity as the start and end of each duplicated span will need to be tracked and considered when training the model.
Deduplication is a very important step and is often used multiple times within the data selection pipeline: first with a simple URL-based filter, then with more complicated hashing- and model-based methods. Deduplication generally relies one of four methods: URLs, hashing, string metrics, or model representations. Bloom filters are often used as state-of-the-art exact deduplication (even though they can remove non-duplicated data) as they are very space efficient. For a more thorough deduplication, approximate matching methods can be used, where MinHashLSH is currently the most commonly used method. [Uncaptioned image] In settings where data efficiency is the primary goal, model-based deduplication has demonstrated good results. However, these methods are relatively new and need further study to demonstrate their benefits in other cases (e.g., when performance is the primary goal). Data Deduplication Summary & Ideas
@@ 3.6 Filtering Toxic and Explicit Content
Practitioners often aim to curate a dataset that is “unbiased and fair”. As such, the aim is for models trained on such data to follow similarly unbiased behavior. Certain popular datasets have also been found to have illegal or extremely inappropriate content, which motivates strict data filters (David, 2023). Towards this purpose, most datasets employ a process to filter out documents that do not conform with such behaviors. Bias and fairness are one of the least well-defined filtering requirements and may be in conflict with different ideologies (Gururangan et al., 2022) and implications (Longpre et al., 2023c). A crucial aspect of this topic is the lack of transparency (Bommasani et al., 2023), requirements, and the downstream effects of applying bespoke bias filtering to pretraining data. Many works solely provide minimal information about the heuristics used for filtering out “biased” data without reporting any ablation on the impact on downstream tasks. A notable exception is the work of Longpre et al. (2023c), which controls and ablates for some of these filters to comprehensively study their effects. They find that using toxicity filters (specifically, Jigsaw’s Perspective API [^13] https://www.perspectiveapi.com/) leads to models that produce less toxic texts, however, the model’s ability to detect toxic content also decreases. As such, one must be careful when using such filters depending on the end-application.
In what follows we summarize the filters related to bias and fairness used in different papers based on the filter types. As mentioned above, such filters are often based on heuristics and the effect of such filtering is not well studied.
@@
Simple filters that aim to remove NSFW content can filter solely based on the source URL, a piece of metadata that is typically stored alongside the text of datasets. Datasets such as RefinedWeb (Penedo et al., 2023) filter out documents coming from URLs that match words from a curated list, as well as from blacklisted domains.
@@
A less common method is to score documents using a model trained only on the undesirable content. Jansen et al. (2022) train a model directly on adult and harmful content and then use the model’s perplexity as a utility function where documents with a high perplexity are assumed to not contain any adult or harmful content.
@@
Finally, Personally Identifiable Information (PII) is another source of concern, which may expose the personal information of individuals. Frequent targeted PIIs are phone numbers, email addresses, IP addresses, and secret keys. A common approach for handling PII is to detect them using regex (Subramani et al., 2023; Elazar et al., 2024; Soldaini et al., 2024; Groeneveld et al., 2024) or classifiers (Allal et al., 2023). Once PII terms are detected, a common practice is to obfuscate them with special tokens.
There is a wide variety of toxic and explicit content that are frequently considered undesirable to have in training or evaluation data. Common methods for filtering out toxic and explicit content include blacklisting certain web domains, blacklisting certain words, using a classifier to detect harmful content, using perplexity to detect harmful content, and removing PII with regular expressions or classifiers. [Uncaptioned image] There have been limited studies on the effects of these filters, warranting further research on the impacts of such filtering to understand how they affect the distribution of data and the downstream models behaviors. [Uncaptioned image] A useful direction for research is how to mitigate any negative impacts of this filtering, or conversely, how can toxic and explicit behaviors be reduced at other stages of model development (e.g., alignment), or possibly even at evaluation time through decoding methods. Toxic and Explicit Content Filtering Summary & Ideas
@@ 3.7 Specialized Selection for Multilingual Models
Many selection methods, such as deduplication, can be reused across languages (Laurençon et al., 2022). However, it is critical to adapt the hyperparameters for the respective language and often manually tune them (Laurençon et al., 2022; Scao, 2023; Scao et al., 2022; Workshop et al., 2022). For example, while heuristics based on technical characters ([0-9{}+/()>]) are still valid, for a length-based filter (discussed in §3.2) one needs to consider that languages like Chinese have much higher information density per character leading to the same content requiring fewer characters than English. Thus, in a filtering pipeline Chinese would require a smaller cut-off value for minimum text length than English (Scao, 2023). Additionally, when using language identification models (e.g., fastText) to detect the language of a document, it is important that the threshold for removal is determined by native speakers, and specific to each language (Laurençon et al., 2022). Not only is the language ID used for determining the language of a document, but it can also be used at the sentence level, where Kudugunta et al. (2023) consider removing sentences if the sentence level language ID and document level ID do not match.
@@
Languages with non-Latin scripts (e.g., Mandarin, Arabic, and Hindi) commonly suffer from incorrect encoding which may require additional filtering or noise correction (Kudugunta et al., 2023). Further, some scripts such as Dongba symbols are not yet supported by Unicode, making the selection of such data especially difficult. Using script-based filters can also serve as language identification for languages that have a one-to-one mapping with a script, such as Japanese Hiragana (§3.1).
@@
Many of the selection methods from other sections in this work can simply be reused when filtering for multilingual data (e.g., deduplication, heuristics on technical characters, data mixing), but some require special parameters for each language (e.g., length-based heuristics). [Uncaptioned image] Particularly for low-resource languages, native speakers must be incorporated into the filtering process to ensure that the data maintains an acceptable quality. Aya (Singh et al., 2024) is a successful example of how this type of inclusion and collaboration can look. Specialized Selection for Multilingual Models Summary & Ideas
@@ 3.8 Data Mixing
Large pretraining corpora often comprise a mixture of data from $k$ domains, where domains are typically defined by provenance. The weighting across domains are specified by the domain weights $\alpha\in\Delta^{k}$, which comprise a $k$-dimensional probability vector. For example, the Pile dataset (Gao et al., 2020) is composed of 24% web data, 9% Wikipedia, 4% GitHub, etc. [^15] The mixture weights depend on the tokenizer. The choice of domain weights can result in significant differences in downstream accuracy (Xie et al., 2023a; Albalak et al., 2023a; Xia et al., 2024a; Fan et al., 2023). Given a fixed training budget, data mixing methods optimize the domain weights to improve training efficiency and model performance.
@@
One challenge that has not been addressed in any of these works is how their weights transfer to new settings. For example, Albalak et al. (2023a) find that the weights determined by Xie et al. (2023a) do not transfer well when using a different tokenizer. It is unclear how to transfer the weights learned under one tokenizer to a new tokenizer without learning the weights from scratch. Furthermore, none of the principled approaches discuss the effects of extrapolating the calculated weights to training on larger quantities of tokens. To exemplify the problem, consider that the weights are calculated by training a model on 50B tokens, and the weights are then used to train a model on 1T tokens. Any domains that saw 1 epoch at 50B tokens will now see 20 epochs at 1T, leading to possible overfitting on smaller domains.
@@
When a pretraining corpus is composed of data from multiple domains, data mixing is an important step of data selection that assigns weights to each domain, determining the final composition of the dataset. Weights can be determined manually, by heuristics, empirically determined, or from principled approaches. Each of these methods have been proven under specific evaluations, but none has proven as a general-purpose best method. [Uncaptioned image] Data mixing is challenging, as there is a natural trade-off when upweighting one domain, the remaining domains must be downweighted. Research on overcoming this trade-off may be difficult, but would be very impactful. One possible direction are methods that consider the interactions (similarities) between domains, which could identify domains that can be downweighted without reducing performance, and those specific domains whose weights cannot be reduced without detriment to performance. [Uncaptioned image] Each of the approaches discussed has different benefits, and combining approaches (e.g., 1/3 heuristic, 1/3 empirically-drive, 1/3 principled) could lead to better general purpose use, but how to best combine them leads to additional complexity. [Uncaptioned image] Although principled approaches are prominent and have lead to state-of-the-art mixing, the weights they determine can also be brittle when used under different settings (e.g., tokenizer, longer training), which are issues that need to be addressed and explored in future research. Data Mixing Summary & Ideas
3.9 Current Best Practices and Landscape For Data Selection
The previous sections have described, in detail, a wide variety of works, methods, and datasets that are relevant to the data selection process. Data selection is an open and active area of research and while the exact details of best practices may change over time, there are some practices and ideas that are of high importance and will likely endure the test of time. In this section we take a step back to look at the high-level of the current landscape of research on data selection, and best practices as they currently stand.
3.9.1 Best Practices
Best practices are actively being developed, and new models are consistently being released with new claims of being state-of-the-art. While there are no blanket statements of what is “best” in all use-cases for language models (because best can only be in reference to a specific evaluation protocol or benchmark, not an actual capability), we describe some trends that have either stood the test of time or have been adopted rapidly and widely. The first trend to notice is that filtering is generally performed starting with the methods that are most efficient and remove broad quantities data, with the later filters being those which are more specialized, but often more costly.
Language filtering. When filtering large quantities of data, an important first step is to ensure that the data is from the desired language, where the most commonly used method is the fastText classifier which has been trained to recognize 157 languages and runs very efficiently.
Heuristic filtering. The next step is usually to remove data using heuristics that are very fast to compute. Examples include removing documents with fewer than 50 or greater than 10,000 words or with fewer than 3 sentences. Examples of commonly used heuristics are those from C4 (Raffel et al., 2020) and MassiveText (Rae et al., 2021). These heuristics can be accessed through the Dolma toolkit (Soldaini et al., 2024) or DataTrove (Penedo et al., 2024).
Data quality. Filtering for “high-quality” data (data which was written by humans and has likely gone through an editing process) is a useful step when processing large quantities of web data, but less likely to be useful for data collected from specific domains (e.g., scientific articles or legal documents). The most common method for quality filtering is to train a classifier that can differentiate between random web data and data from a known high-quality domain. There are no commonly used open-source models, but the methods have been described within previous works (Brown et al., 2020b). A common approach when using quality filters is to use a stochastic filter, which will allow some data with a low quality score to pass through. This can help to reduce some of the biases known to exist in quality filters.
Domain-specific filtering. Methods for domain-specific selection are quite similar to the methods for data quality, but only required when training a model for a specific target domain. The goal of these methods will be to find data within a large general purpose corpus which is most similar to a target domain. The current best methods, such as DSIR (Xie et al., 2023b), train two models: one in-domain and one general domain model. To determine whether a data point is similar to the target domain, the difference of probabilities under each model is calculated.
Deduplication. Deduplication has proven to be a beneficial step in every pretraining data selection pipeline. There are multiple methods of deduplication that vary from very simple and broad deduplication methods which only consider metadata to very fine-grained deduplication on individual sentences. When using web-text, a simple and cheap first step is to deduplicate data based on URLs. Next, hash-based methods such as MinHash (Broder, 1997) (approximate matching) and Bloom filters (exact matching) have been used with great success. The previously mentioned methods are fairly cheap, but only consider word or token matching, and do not consider semantic matching. Model-based methods such as SemDeDup (Abbas et al., 2023) have demonstrated success for deduplicating semantically similar data, but are much more computationally expensive and should generally be applied later in the filtering pipeline after cheaper filtering methods.
Filtering toxic and explicit content. Removing toxic and explicit content from web text is generally performed through multiple steps to handle different types of content. NSFW content can be removed, in part, by filtering out URLs that match NSFW word lists and from blacklisted domains. Beyond just searching for bad words in the URL, it is also common to filter out documents that contain the same list of NSFW words. Removing personally identifiable information (PII) is another very important step and is commonly found and removed using regex filters, a good example of which is Dolma (Soldaini et al., 2024).
Selection for multilingual models. Generally speaking, most methods from the other steps will work the same for multilingual models, with the exception that some hyperparameters may need to be adjusted to work for each language. An important best practice for creating multilingual models is to incorporate native speakers into the decision making process to ensure that data maintains an acceptable quality.
Data mixing. Data mixing is only used when the training dataset is composed of data from multiple clearly defined domains or sources (e.g., books or web text). As of writing, there are no clear best practices for automated methods for data mixing as this is still a very young area of research. Methods of determining domain weights automatically tend to use loss-per-domain as part of a signal as to the importance/difficulty of each domain. Starting points for research are DoReMi (Xie et al., 2023a) and ODM (Albalak et al., 2023a).
3.9.2 Current Landscape of Datasets For Data Selection Research
The most popular datasets used for data selection until now have been RedPajama (Skill-it (Chen et al., 2023c), SlimPajama (Shen et al., 2024), data mixing laws (Ye et al., 2024)), The Pile (ODM (Albalak et al., 2023a), DSIR (Xie et al., 2023b), DoReMi (Xie et al., 2023a), Selection for encoders (Feng et al., 2022)), and C4 (DsDm (Engstrom et al., 2024)), each with pros and cons. C4 (Raffel et al., 2020) (750Gb of data) is the oldest and contains only webcrawled data. C4 generally leads to models with lower benchmark scores as the other 2 datasets, but this makes it an easy first testing ground for data selection methods since there is increased room for improvement. The Pile (Gao et al., 2020) (800Gb of data) dataset contains slightly more data than C4, but only ~28% of the total dataset comes from web scrapes and the remaining 72% comes from a wide variety of domains including books, scientific articles, and Github. This combination of well-defined sources makes the Pile a good candidate for research on data mixing. Lastly, RedPajama (1T tokens) is the newest and the largest of the commonly used datasets, and was originally intended to be an open-source re-creation of the dataset used to train Llama-2. Recently, RedPajama-2 (30T tokens) was released which may be an even better resource for testing data selection. RedPajama-2 comes with precomputed quality signals including those implemented for C4 and RefinedWeb, allowing for experimentation on selection mechanisms under controlled settings. Another beneficial aspect of RedPajama-2 is that it contains a huge amount of data, allowing researchers to filter the data down to various scales, ranging from billions to trillions of tokens.